what do you understand by the entropy of a substance ? what are its units ? write the general expression for the entropy change of a process.
1
Expert's answer
2019-10-28T07:49:17-0400

Solution.

1.The entropy of matter (from thermodynamics) is a state function that is a measure of the disorderliness of a thermodynamic system.

2.The unit of entropy is J/(Mol*K).

3.The change in the entropy of a chemical reaction (process) is expressed by the following formula:

"\Delta S = \sum (\nu S0(product)) - \sum(\nu S0(star.))"

Answer:

1.The entropy of matter (from thermodynamics) is a state function that is a measure of the disorderliness of a thermodynamic system.

2.The unit of entropy is J/(Mol*K).

3.The change in the entropy of a chemical reaction (process) is expressed by the following formula:

"\Delta S = \sum (\nu S0(product)) - \sum(\nu S0(star.))"